when i was 2yrs old my sister was half my age .now i'm 100yrs old how old is my sister now?
Answer by thejackal(72)

when I was 2, S was 2/2 = 1
so the age difference between I and Sister is 1 year
so if I is 100 then S = 100-1 = 99